# SegMASt3R: Geometry Grounded Segment Matching

### v1

@@ 6 Conclusions
We proposed SegMASt3R, a simple method to re-purpose an existing 3D foundation model MASt3R for image segment matching. Our proposed method achieves excellent results on ScanNet++ and Replica with a simple pipeline and a minimal amount of training. It especially excels on wide-baseline segment matching between image pairs. In addition, we show that SegMASt3R has practical applicability by evaluating it’s performance on Instance Mapping and Image goal navigation as downstream tasks. For both applications, it outperforms compared to baselines.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
@@ Appendix A Technical Appendices and Supplementary Material
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We re-purpose a 3d foundation model and show how it can be used for segment matching. We show experiments showcasing the efficacy of our proposed method and down-stream applications justifying its use in real-world robotic settings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.

### v2

@@ 6 Conclusion
We proposed SegMASt3R, a simple method to re-purpose an existing 3D foundation model MASt3R for image segment matching. Our proposed method achieves excellent results on ScanNet++ and Replica with a simple pipeline and a minimal amount of training. It especially excels on wide-baseline segment matching between image pairs. In addition, we show that SegMASt3R has practical applicability by evaluating it’s performance on the downstream tasks of 3D instance mapping and object-relative topological navigation, where we significantly outperform the corresponding baselines. Overall, this work attempts to establish segment matching as a core computer vision capability, which will enable even more downstream applications in future alongside the advances in image segmentation and data association.
Supplementary Material
Appendix A Cross-Dataset Generalization on Replica
To evaluate cross-dataset generalization, we train our model on ScanNet++[50] and directly evaluate it on the Replica dataset[42], without any fine-tuning. Table 8 reports a detailed comparison against competitive baselines, highlighting the robustness of our approach under distribution shift.
Type Method 0°–45° 45°–90° 90°–135° 135°–180°
AUPRC R@1 R@5 AUPRC R@1 R@5 AUPRC R@1 R@5 AUPRC R@1 R@5
Local Feature Matching (LFM) SP-LG [11, 29] 64.69 67.89 72.3 46.89 51.59 56.1 21.86 28.74 32.83 21.82 28.72 30.9
GiM-DKM [39, 13] 75.46 81.38 83.51 67.26 74.08 76.8 43.94 55.8 61.41 35.37 49.41 53.86
[…]
MASt3R [24] 52.2 57.5 81.2 39.1 51.0 78.6 23.6 45.9 77.2 17.2 43.8 75.7
[…]
Table 8: Performance of additional baselines across pose‐bins on Replica[42]. Blue cells mark the best scores; Orange cells mark the second‐best.
Appendix B Local-Feature Matching (LFM) Baseline Setup
Algorithm 1 Segment Correspondence via Keypoint Voting
1:Images $I_{0},I_{1}$; segment masks $\mathcal{M}_{0}\in\{0,1\}^{M\times H\times W}$, $\mathcal{M}_{1}\in\{0,1\}^{N\times H\times W}$; keypoint matcher $\mathcal{A}$; max matches $K$
2:Segment assignment vector $\hat{\mathcal{C}}\in\mathbb{N}^{M}$ (with -1 indicating no confident match)
[…]
4:Initialize vote matrix $V\in\mathbb{Z}^{M\times N}\leftarrow 0$
5:for $i=1$ to $K$ do
6:  Find $m_{i}$ such that $\mathcal{M}_{0}[m_{i},y_{i}^{0},x_{i}^{0}]=1$
7:  Find $n_{i}$ such that $\mathcal{M}_{1}[n_{i},y_{i}^{1},x_{i}^{1}]=1$
8:  if $m_{i}$ and $n_{i}$ are valid then
9:   $V[m_{i},n_{i}]\leftarrow V[m_{i},n_{i}]+1$
10:  end if
11:end for
12:for $m=1$ to $M$ do
13:  if $\sum_{n}V[m,n]=0$ then
14:   $\hat{\mathcal{C}}[m]\leftarrow-1$ $\triangleright$ No confident match
15:  else
16:   $\hat{\mathcal{C}}[m]\leftarrow\arg\max_{n}V[m,n]$
17:  end if
18:end for
19:return $\hat{\mathcal{C}}$
A range of local feature matchers (LFMs) exist at varying densities—SuperPoint [11] (sparse), LoFTR [43] (semi-dense), and RoMa or MASt3R [14, 24] (dense). We leverage the EarthMatch (IMM) [5] toolkit, which provides a unified and modular interface for applying a wide range of local feature matchers. This abstraction significantly simplifies the integration of keypoint-based methods into our pipeline, allowing us to systematically evaluate segment correspondence performance across multiple matcher densities.
We use a voting-based aggregation scheme (Algorithm 1) to derive segment-level correspondences via LFMs. Given two images and their respective binary segment masks $\mathcal{M}_{0}\in\{0,1\}^{M\times H\times W}$ and $\mathcal{M}_{1}\in\{0,1\}^{N\times H\times W}$, a chosen LFM $\mathcal{A}$ produces $K$ matched keypoint pairs. Each matched pair votes for the segment it belongs to in the source and target masks, incrementing a vote matrix $V\in\mathbb{Z}^{M\times N}$.
Finally, segment-level correspondences are assigned by selecting, for each source segment $m$, the target segment $n$ with the highest vote count. Segments with no valid votes are marked as unmatched (assigned label -1). This voting mechanism provides a simple yet effective bridge between pixel-level keypoint matches and object-level segment associations.
Appendix C Additional Qualitative Results on ScanNet++ Dataset
We present additional qualitative results (Figure 5) on the ScanNet++ dataset[50], highlighting our method’s robustness to perceptual instance aliasing and wide-baseline viewpoint changes. These examples further demonstrate accurate segment correspondences despite challenging geometric and appearance ambiguities.
Figure 5: More examples comparing proposed method against DINOv2 [32] and SAM2 [37] for segment-matching on the ScanNet++ Dataset [50] under Wide-baseline conditions [135–180° viewpoint change]. Both baselines tend to incorrectly assign segment correspondences given an opposite viewing direction-the tables (pink and green) in (a) as well as the chairs (red and blue) in (c). Another failure mode SAM2 specifically exhibits is the inability to propagate masks in challenging view-point change settings as seen for the wheelchair (green) in (b). In contrast, the proposed method demonstrates accurate segment matching, attributed to its conditioning on 3D-aware priors.
Note: Images have been greyed out to improve the visibility of the segments in question.
Appendix D Details on Instance-Mapping
Method office0 office1 office2 office3 office4 room0 room1 room2
ConceptGraphs (MobileSAM Masks) [19] 11.84 / 28.43 20.31 / 43.79 8.63 / 22.82 8.07 / 22.83 9.46 / 24.73 12.23 / 34.34 5.83 / 12.96 7.83 / 23.82
ConceptGraphs (GT Masks) [19] 43.53 / 69.68 22.48 / 40.71 43.46 / 60.69 32.06 / 53.44 39.63 / 68.22 44.89 / 69.64 17.96 / 36.53 25.93 / 43.63
[…]
By default, ConceptGraphs [19] is evaluated on the scanned RGB-D trajectories of the Replica [42] as provided by Nice-SLAM [54]. However, this version does not include ground-truth instance masks and instead relies on MobileSAM [51] for mask generation. We first evaluate the method using this setup, which corresponds to the first row of Table 9. To assess performance with ground-truth instance segmentation, we use the pre-rendered Replica RGB-D sequence released by Semantic-NeRF [53], which provides 900 RGB-D frames per scene along with ground-truth instance masks and camera poses. We subsample each sequence with a stride of 5 and run ConceptGraphs on a total of 180 frames per scene. This evaluation setting corresponds to the second row of Table 9. We observe that the trend remains consistent, with SegMASt3R continuing to outperform ConceptGraphs on the instance mapping task.
D.1 Evaluation Methodology for Instance Mapping
We assess the accuracy of instance mapping by how well individual object geometries are preserved, i.e. quality of the instance boundaries and whether they are multi-view consistent. We follow the class-agnostic evaluation protocol as done by OpenYOLO3D [6], which benchmarks generic 3D object instances in the given scene. Formally, given a predicted 3D instance-map point-cloud and the corresponding ground-truth, we compute the Average Precision (AP) over a range of Intersection-over-Union (IoU) thresholds. In particular, we report:
  • AP@0.50 reports precision when IoU = 0.50, measuring correctness at a moderate overlap.
  • AP is the mean precision over IoU thresholds from 0.50 to 0.95 in steps of 0.05, capturing both coarse and fine alignment.
(a) Ground Truth
(b) ConceptGraphs
(c) Ours
Figure 6: Qualitative comparison of instance maps. Each color corresponds to a different object instance. The ground truth (left) provides the closest reference RGB view of the scene from the RGB sequence. ConceptGraphs (middle) tends to over-segment objects, assigning multiple instance IDs to the same object, resulting in fragmented coloring (e.g. brown duvet covered by green and yellow labels). Our method (right) produces cleaner and more consistent instance groupings.
D.2 ConceptGraphs for Instance Mapping
We use Concept-Nodes, a lightweight implementation of ConceptGraphs [19]. We generate an instance-level 3D map by aggregating detections over time and associating them with previously observed objects in a global map via semantic and geometric matching.
D.2.1 Object-Centric 3-D Representation
Given a stream of RGB-D observations $\mathcal{I}=\{I_{1},\dots,I_{T}\}$, ConceptGraphs incrementally builds a 3-D scene graph $\mathcal{M}_{t}=(\mathcal{O}_{t})$. Each node $o_{j}\in\mathcal{O}_{t}$ is represented by a point cloud $\mathbf{P}_{o_{j}}$ and a unit-normalised semantic descriptor $\mathbf{f}_{o_{j}}$. At time $t$ the incoming frame $I_{t}=\langle I_{t}^{\mathrm{rgb}},I_{t}^{\mathrm{depth}},\boldsymbol{\theta}_{t}\rangle$ (colour, depth, pose) is fused into the map by either updating an existing object or instantiating a new one.
D.2.2 Class-Agnostic 2-D Segmentation
For each frame we extract $M$ class-agnostic masks $\{\mathcal{M}_{t,i}\}_{i=1}^{M}=\mathrm{Seg}(I_{t}^{\mathrm{rgb}})$ using YOLO-World [9] and MobileSAM [51] (or ground truth, when available). Every mask is fed to a CLIP [35] encoder to obtain a visual descriptor $\mathbf{f}_{t,i}=\mathrm{SE}(I_{t}^{\mathrm{rgb}},\mathcal{M}_{t,i})$. The masked RGB-D region is back-projected and transformed to the global frame, yielding a point cloud $\mathbf{P}_{t,i}$ paired with $\mathbf{f}_{t,i}$.
D.2.3 Object Association
For every newly detected object candidate $\langle\mathbf{P}_{t,i},\mathbf{f}_{t,i}\rangle$ we measure its similarity to each map object $\langle\mathbf{P}_{j},\mathbf{f}_{j}\rangle$ that exhibits spatial overlap. Geometric consistency is captured by the nearest-neighbour ratio
$s_{\text{geo}}(i,j)=\operatorname{NNRatio}(\mathbf{P}_{t,i},\mathbf{P}_{j}),$
defined as the fraction of points in the candidate cloud whose closest point in the reference cloud lies within the Euclidean tolerance $\delta_{\text{nn}}$. Appearance similarity is obtained from the cosine similarity of the CLIP descriptors, rescaled from [1,1][-1,1] to $[0,1]$,
$s_{\text{sem}}(i,j)=\tfrac{1}{2}\bigl(\mathbf{f}_{t,i}^{\!\top}\mathbf{f}_{j}\bigr)+\tfrac{1}{2}.$
The two scores are linearly blended into a fused similarity
$s(i,j)=\alpha\,s_{\text{sem}}(i,j)+(1-\alpha)\,s_{\text{geo}}(i,j),\qquad\alpha\!\in\![0.1,0.5],$
where the weight $\alpha$ balances semantic and geometric evidence and is tuned per scene. Greedy matching assigns each detection to the map object with the highest fused similarity; if this peak score falls below the threshold $\delta_{\text{sim}}\in[0.90,0.96]$, the detection is instead used to instantiate a new object instance. In practice, we sweep and select $\delta_{\text{sim}}$ in the range $0.90\dots 0.96$ to accommodate scene-specific variability in texture and clutter.
D.2.4 Object Fusion
Once a detection $i$ has been matched to map object $o_{j}$, its semantic and geometric information is merged into the map. Let $n_{j}$ denote the number of detections previously fused into $o_{j}$ and $\delta_{\text{voxel}}$ the voxel size used for down-sampling.
Semantic update.
We maintain a running average of CLIP descriptors so that each new observation contributes equally:
$\mathbf{f}_{j}\;\leftarrow\;\frac{n_{j}\,\mathbf{f}_{j}+\mathbf{f}_{t,i}}{n_{j}+1},\qquad n_{j}\;\leftarrow\;n_{j}+1.$
Geometric update.
The point cloud of the object is augmented and then compacted to prevent redundancy:
$\mathbf{P}_{j}\;\leftarrow\;\operatorname{Downsample}\!\bigl(\mathbf{P}_{j}\;\cup\;\mathbf{P}_{t,i},\delta_{\text{voxel}}\bigr),$
where $\operatorname{Downsample}(\cdot,\delta_{\text{voxel}})$ groups points whose pairwise distances are below the voxel size ($\delta_{\text{voxel}}\!=\!1\text{\,cm}$ in all experiments) and replaces each group by its centroid.
D.2.5 SegMASt3R for Instance Mapping
We use SegMASt3R to directly establish object-level correspondences across image pairs, based on both mask-level appearance features and geometric consistency. Unlike Concept-Graphs, which incrementally associates detections with a growing map, our method operates pairwise and enforces geometric alignment through 3D checks.
Pairwise Mask Matching.
Given two RGB-D frames $I_{s}=\langle I_{s}^{\text{rgb}},I_{s}^{\text{depth}},\theta_{s}\rangle$ and $I_{t}=\langle I_{t}^{\text{rgb}},I_{t}^{\text{depth}},\theta_{t}\rangle$ and their respective sets of ground-truth instance masks $\{\mathbf{m}_{s,i}\}_{i=1}^{M_{s}}$ and $\{\mathbf{m}_{t,j}\}_{j=1}^{M_{t}}$, we extract visual descriptors for each mask using our method, SegMASt3R:
$\mathbf{f}_{s,i},\mathbf{f}_{t,j}=\textsc{SegMASt3R{}}(I_{s}^{\text{rgb}},I_{t}^{\text{rgb}},\mathbf{m}_{s,i},\mathbf{m}_{t,j})$
where $\mathbf{f}_{s,i},\mathbf{f}_{t,j}\in\mathbb{R}^{24}$ are the mask feature vectors. These features are used to construct a pairwise similarity matrix, and a soft assignment between masks is computed using the Sinkhorn algorithm. Tentative correspondences are obtained by row-wise $\arg\max$ over the resulting transport matrix:
$j^{*}=\arg\max_{j}\hat{M}_{i,j},$
where $\hat{M}\in\mathbb{R}^{M_{s}\times M_{t}}$ is the normalized assignment matrix. Matches assigned to the dustbin (i.e., $j^{*}=M_{t}+1$) are discarded. The remaining matches are considered for geometric consistency.
Geometric Consistency.
For each matched pair $(\mathbf{m}_{s,i},\mathbf{m}_{t,j^{*}})$, we project both masks into 3D using the depth maps and camera poses to obtain their corresponding point clouds, $\mathbf{p}_{s,i}$ and $\mathbf{p}_{t,j^{*}}$. To evaluate geometric consistency, we compute the geometric similarity between the two point-clouds in the same way as done by Concept-Graphs. A match is accepted if $\phi_{\text{geo}}(i,j^{*})>0.5$. All matches failing this geometric filter are rejected.
Evaluation on Replica.
We use the RGB-D frames from Semantic-NeRF [53] and subsample each 900-frame sequence with a stride of 20, resulting in 45 frames per scene. We perform pairwise matching across all possible frame pairs (990 in total) and aggregate the correspondences to construct the final instance map. The corresponding results are reported in the third row of Table 9.
Appendix E Additional Details on Object-Relative Topological Navigation
Background: We used RoboHop [17] for benchmarking on the topological navigation task. RoboHop uses a segment-based topological map to enable open-world visual navigation. It extracts semantically-meaningful image segments using SAM [22]. These segments become the nodes of a topological graph, where edges are formed both intra-image–linking segments within the same image based on pixel-level proximity–and inter-image–connecting corresponding segments across consecutive frames. For given query image segments, segment matching is performed using SuperPoint+LightGlue to retrieve matching segment nodes from the map, each of which has a precomputed path length to the given goal. Using these path lengths along with the pixel centers of the image segments, yaw ($\psi$) is computed as below [34, 17]:
$\psi=\frac{K}{W}\sum_{i}w_{i}(x_{i}-c);\quad w_{i}=\frac{e^{\tau l_{i}}}{\sum_{i}{e^{\tau p_{i}}}}$ (5)
where $c$ is the image center, $x_{i}$ is the segment center, $p_{i}$ is the path length, $\tau$ is the temperature parameter (set to $5$), $w_{i}$ is the softmax weight per query segment, $W$ is the image width, and $K$ is the proportional gain (set to $0.4$). For forward translation, a fixed velocity of $0.5m/s$ is used. We obtained the source code from the original authors of RoboHop [17], and followed [34] to evaluate using HM3Dv0.2 [36]. This benchmark data is based on the validation set of the InstanceImageNav (IIN) challenge [23], where $108$ episodes ($3$ each from the $36$ unique environments) are used to define the start position, object goal, and map trajectory. A topological graph is constructed a priori from the map trajectory, which is then used during the execution phase for localization, planning, and control.
Figure 7: Different components of the image based goal navigation video.
Qualitative Visualizations: In our experiments, we replaced RoboHop’s segment matching during localization with SegMASt3R, and compared vanilla RoboHop with this version. In this supplementary material, we include navigation videos that qualitatively compare the two methods. The left half of the video corresponds to vanilla RoboHop, while the right half corresponds to our SegMASt3R-enhanced version. Here, we provide an overview of the visualization panel of the video. At the top, we show an overhead view of the simulator, where the map trajectory is displayed with a color gradient starting from the blue circle and ending at the green diamond (goal object). The agent is initialized $5m$ away from the goal along the map trajectory, and its current state is displayed with a white circle. In the middle, we show the image segments which are color-coded based on the path lengths (normalized per image): yellow represents desirable segments (high weight $w_{i}$) that lead to the goal and blue represents the undesirable segments. The arrows on the segments indicate yaw direction ($\psi$), and the length of the arrow reflects its horizontal pixel offset from the image center ($(x_{i}-c)$). At the bottom, we show the egocentric RGB view from the agent’s current state.
We compare our method against the vanilla RoboHop baseline on four HM3D episodes via side-by-side video clips, emphasizing two key aspects:
@@ Appendix E Additional Details on Object-Relative Topological Navigation
    Success vs. Failure: nav_bed_20.mp4 and nav_chair_55.mp4 demonstrate that, with our segment-matching module, the agent successfully reaches the goal, whereas with the vanilla RoboHop method it fails to reach the goal.